the sum of two numbers is 36. the first number is three more than twice the second number. find the larger number
Answer by mananth(16949)

let one number be x
The other number will be 36-x
x= 3+2(36-x)
x= 3+72-2x
3x=75
x=25
36-25=11
the numbers are 11 & 25